OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 4 - 3x. -3/8 -8/3 2

OpenStudy (anonymous):

get it into the form y = mx + b

OpenStudy (anonymous):

8y=-3x+4

OpenStudy (anonymous):

isolate the y..do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y=-3/8x+4

OpenStudy (anonymous):

if you divide the equation by 8..you will isolate y y =-3x/8 + 4/8 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now in this form what is the coeffiecient of x?

OpenStudy (anonymous):

-3/8

OpenStudy (anonymous):

so remember ..in the form y=mx+b m is always the slope so what is the slope now?

OpenStudy (anonymous):

m=-3/8

OpenStudy (anonymous):

yes..good..so thats the slope..got it?